A rm's total cost function is given as T C = 4 + 8Q − Q 2 , and total revenue as T R = 50Q − 4Q 2 , where Q is the number of units produced and sold (in thousands). How many units should be produced to maximize prot? 


1
Expert's answer
2021-07-12T13:12:00-0400

Solution.


"TC=4+8Q-Q^2 \newline
TR=50Q-4Q^2"


Profit = TR - TC = 50Q - 4Q2 - (4 + 8Q - Q2) = 50Q - 4Q2 - 4 - 8Q + Q2 = -4 + 42Q - 3Q2

Profit has maximum than "\frac{d Profit}{d Q} = 0,\newline
42 - 3*2*Q=42-6Q=0."From here Q = 7 (thousands)

Answer: 7